OpenStudy (anonymous):

Identify the 27th term of an arithmetic sequence where a1 = 38 and a17 = -74. -20.5 -151 -22.75 -144

OpenStudy (campbell_st):

ok to find a term in an arithmetic sequence use \[a_{n} = a + (n - 1) \times d\] the information in your question can be used to find the common difference , d \[-74 = 38 + (17 -1) \times d.......or..... -74 = 38 + 16d\] solve for d. then substitute n = 27, a = 38 and the value of d into the formula above to find the 27th term

OpenStudy (anonymous):

I got 227 but none of those are the answers?

OpenStudy (campbell_st):

well I think you should know the common difference is negative -112 = 16d d = -7 so the 27th term \[a_{27} = 38 + (27 -1) \times (-7)\]

OpenStudy (anonymous):

Thank you!